Circular No. 6100

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

WGACAT N. E. White, Laboratory for High Energy Astrophysics, Goddard Space

Flight Center (LHEA/GSFC); P. Giommi, European Space Information System, European Space Research Institute; and L. Angelini, LHEA/GSFC, report that WGACAT, a ROSAT catalogue of point sources, is now available through HEASARC, the High Energy Astrophysics Science Archive Research Center online service. This catalogue, generated using the ROSAT PSPC pointed data publicly available up to Sept. 1994, contains 50 000 detections and more than 45 600 individual sources; a total of 2624 fields were searched. WGACAT is an independent research effort aimed at releasing as quickly as possible a list of sources detected by ROSAT in its pointed phase to identify the detected sources. Optical identification of these xray sources is encouraged. WGACAT can be accessed via HEASARC by logging into legacy.gsfc.nasa.gov with the username xray. Type browse wgacat to access the catalogue. WGACAT can also be accessed via the anonymous ftp account on the same machine under the directory rosat/wgacat/files and via the HEASARC WWW Browse interface un-

der http://heasarc.gsfc.nasa.gov/StarTrax/Browse.html by clicking
on ROSAT and then selecting WGACAT. Send mail to wgacat@athena.gsfc.

nasa.gov if you have any questions regarding this catalogue.

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Ephemeris extension (viz. IAUC 6070) from the orbit on MPC 23956:

1994/95 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 25 20^{h}58.^{m}26 −46^{o}08.^{′}4 1.645 1.947 91.^{o}7 30.^{o}7 12.0
Nov. 4 21 05.57 −47 06.6 1.916 2.060 83.8 28.6 12.5
14 21 15.24 −47 36.9 2.178 2.173 76.6 26.3 13.1
24 21 26.61 −47 49.7 2.431 2.287 69.9 23.9 13.5
Dec. 4 21 39.22 −47 50.7 2.671 2.401 63.6 21.6 13.9
14 21 52.77 −47 43.9 2.896 2.515 57.9 19.4 14.3
24 22 07.00 −47 32.3 3.106 2.628 52.6 17.3 14.7
Jan. 3 22 21.73 −47 17.9 3.299 2.740 48.0 15.5 15.0
13 22 36.84 −47 02.5 3.474 2.852 44.2 13.9 15.3
23 22 52.20 −46 47.7 3.631 2.963 41.3 12.7 15.5
Feb. 2 23 07.76 −46 34.8 3.769 3.074 39.5 11.8 15.8

Total visual magnitude estimate by A. Hale, near Orogrande, NM (0.20-m reflector): Sept. 28.18 UT, 10.7.

1994 November 3 Brian G. Marsden